house ramen's current capital structure

House Ramen’s current capital structure is 70% equity, 25% debt, and 5% preferred stock. This is considered optimal. House Ramen is considering a $50 million capital budgeting project. During the coming year they expected to have $15 million of retained earnings available to finance this capital budgeting project. The marginal tax rate is 40.00%.

·Long-term debt can be raised at a pretax interest rate of 7.00%

·Preferred stock can be sold at a $25 price with a $2 annual dividend. Flotation or issuance costs will be $3 per preferred share.

·Common stock can be sold at a $20 price. The common dividend is expected to be $3.00 next year. Dividends have been growing at an annual compound rate of 4% annually and are expected to continue growing at that rate into the foreseeable future. Flotation or issuance costs will be $4 per common share.

Calculate House Ramen’s weighted average cost of capital.

Weighting of the equity component can be divided into internal equity and external equity. While the total is 70%, the amount for internal equity is _____% and for external equity is _____%.

after-tax cost of debt is _____ %, cost of preferred stock is _____%, cost of internal equity is _____% and cost of external equity is _____%. 

Therefore, the weighted average cost of capital is _____%. 
Answer #1

1.

30%


2.

40%


3.

=7%*(1-40%)

=4.200%


4.

=2/(25-3)

=9.091%


5.

=3/20+4%=19.00%


6.

=3/(20-4)+4%=22.75%


7.

=25%*7%*(1-40%)+30%*(3/20+4%)+40%*(3/(20-4)+4%)+5%*(2/(25-3))

=16.305%


answered by: aux